OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4+ 9x^2

OpenStudy (aum):

Factor x^2 out. What do you get?

OpenStudy (anonymous):

im clueless where do I start

OpenStudy (aum):

Start with factoring out x^2. What do you get when you factor out x^2?

OpenStudy (anonymous):

x^2(x^2)

OpenStudy (anonymous):

@aum

OpenStudy (aum):

f(x) = x^4+ 9x^2 = x^2(x^2 + 9) They want linear factor and therefore the highest power of each factor should be 1. f(x) = x^2(x^2 + 9) = x * x * (x^2 + 9) = x * x * (x + 3i) * (x - 3i)